A school is making a rectanglular flag. The width of the flag is 25 inches long. The ratio of the length-to-width is 19:10. What is the length?
Answer by checkley77(12844)

19:10 X:25
10X=25*19
10X=475
X=47.5 IN. FOR THE LENGTH.
PROOF
19/10=47.5/25
475=475